A random sample of size 50 is drawn from binomial distribution with parameters n=100 and p=0.3. what is the probability that the sample mean is



a)larger than 3.9



b) between 4.1 and 4.4



c) smaller than 4.0

1
Expert's answer
2022-02-17T03:34:00-0500

Let "X" be a binomial rv based on "n" trials with success probability "p." Then if the

binomial probability histogram is not too skewed, "X" has approximately a

normal distribution with "\mu=np" and "\sigma=\sqrt{npq}."

In practice, the approximation is adequate provided that both "np\ge 10" and "nq\ge 10."

Given "n=100, p=0.3, q=1-p=1-0.3=0.7."

"np=100(0.3)=30\ge 10"

"nq=100(0.7)=70\ge 10"

We can use normal approximation for binomial distribution with "\mu=np=100(0.3)=30, \sigma=\sqrt{npq}=\sqrt{100(0.3)(0.7)}=\sqrt{21}"

Let "\bar{X}=" the sample mean: "\bar{X}\sim N(\mu, \sigma^2/n_1)"

Given "n_1=50."

a)


"P(\bar{X}>3.9)=1-P(Z\le\dfrac{3.9-30}{\sqrt{21}/\sqrt{50}})"

"\approx1-P(Z\le-40.2732)\approx1"


b)


"P(4.1<\bar{X}<4.4)=P(Z<\dfrac{4.4-30}{\sqrt{21}/\sqrt{50}})"

"-P(Z\le\dfrac{4.1-30}{\sqrt{21}/\sqrt{50}})"




"\approx P(Z<-39.5017)-P(Z\le-39.9646)\approx0"


c)


"P(\bar{X}<4.0)=P(Z<\dfrac{4.0-30}{\sqrt{21}/\sqrt{50}})"

"\approx P(Z<-40.1189)\approx0"